OpenStudy (anonymous):

Inversely, Direct or Neither (Variations) . y/x = 3 -5 = xy y = 8x

OpenStudy (anonymous):

I say last was is direct and 1st two are neither because 1st one is set to y/x not y or x and the 2nd one is xy but not to in respect to x or y.

OpenStudy (anonymous):

Last one* not last was

OpenStudy (anonymous):

This is what they look like: 1) \[\frac{ y}{x }=3\] 2) \[-5=xy\] 3)\[y=8x\]

OpenStudy (anonymous):

Yeah I know. #1 and #2 are neither and #3 is direct

OpenStudy (mertsj):

It's easier to tell if you make them say y =

OpenStudy (anonymous):

Look two posts up, I did.

OpenStudy (mertsj):

\[\frac{y}{x}\times x=3x\] y=3x

OpenStudy (mertsj):

Clearly direct variation.

OpenStudy (anonymous):

So that's for the first one?

OpenStudy (anonymous):

Oh no, the third one I mean?

OpenStudy (mertsj):

Where did you solve them for y? I can't find it.

OpenStudy (anonymous):

That's how the equation is shown. It's not being solved, it's just asking to state what kind of variation it is.

OpenStudy (anonymous):

Yeah I was about to say the question didn't say it needed to be solved

OpenStudy (mertsj):

Number 1: \[\frac{y}{x}=3\] \[\frac{y}{x}\times x=3\times x\] \[y=3x\]

OpenStudy (anonymous):

So that's direct?

OpenStudy (mertsj):

Number 2 \[-5=xy\] \[xy=-5\] \[\frac{xy}{x}=\frac{-5}{x}\] \[y=\frac{-5}{x}\]

OpenStudy (mertsj):

Number 3 \[y=8x\]

OpenStudy (mertsj):

Now which one(s) are of the form y =kx? Those are direct variation. Which one(s) are of the form : \[y=\frac{k}{x}\] Those are indirect variation.

OpenStudy (anonymous):

Do you mean inverse variation? & Would it be #3?

OpenStudy (mertsj):

Indirect and inverse are the same thing.

OpenStudy (anonymous):

Oh okay , haha sorry. So, #1 is indirect, #2 is ? , #3 is direct?